# Gödel Agent: A Self-Referential Framework for Agents Recursively Self-Improvement

### v2

@@ 4.3 Experimental Results and Analysis
The experimental results on the four datasets are shown in Table 1. Under the same experimental settings, Gödel Agent achieves either optimal or comparable results to Meta Agent Search across all tasks. Notably, in the mathematics task MGSM, Gödel Agent outperforms the baseline by 11%. This suggests that reasoning tasks offer greater room for improvement for Gödel Agent, while in the knowledge-based QA dataset, it only slightly surpasses baselines. In contrast to Meta Agent Search, which relies on manually designed algorithmic modules to search, Gödel Agent demonstrates greater flexibility. It requires only a simple initial policy, such as CoT, with all other components being autonomously generated. Moreover, through interaction with the environment, Gödel Agent gradually adapts and independently devises effective methods for the current task. The final policies generated by Gödel Agent for four tasks are shown in Appendix C.1. Additionally, our method converges faster, with the required number of iterations and computational cost across different tasks compared to the Meta Agent shown in Appendix D.
We also conduct experiments without restrictions, where Gödel Agent significantly outperforms all baselines. Upon further analysis, we find that this is primarily due to the agent’s spontaneous requests for assistance from more powerful models such as GPT-4o in some tasks. Therefore, Gödel Agent is particularly well-suited for open-ended scenarios, where it can employ various strategies to enhance performance.
@@ 5 Analysis
To further explore how Gödel Agent self-improves, as well as the efficiency of self-improvement and the factors that influence it, we first evaluate the tool usage ratio on the MGSM dataset and conduct an ablation study on the initial tools. In addition, to analyze the robustness of Gödel Agent’s self-improvement capabilities, we also collect statistics on factors such as the reasons for the agent’s termination. Finally, we perform a case study of initial policies and optimization processes on the classic Game of 24.
5.1 Analysis of Initial Tools
We record the number of different actions taken in the experiments. As shown in Figure 3, we can see that Gödel Agent interacts with its environment frequently, analyzing and modifying its own logic in the process. Additionally, error handling plays a crucial role.
As discussed in Section 3.2, Gödel Agent is initially provided with four additional tools to accelerate convergence and reduce optimization difficulty: 1) thinking before acting, 2) error handling, 3) code running, and 4) LLM calling. To analyze their impact, an ablation study is conducted, and the results are shown in Table 2.
Table 2: Ablation study on initial tool configuration.
Different Actions MGSM
Gödel Agent 64.2
w/o thinking 50.8
w/o error handling 49.4
w/o code running 57.1
w/o LLM calling 60.4
The study reveals that the “thinking before acting” tool significantly influences the results, as much of Gödel Agent ’s optimization effectiveness stems from pre-action planning and reasoning. Additionally, error handling is crucial for recursive improvement, as LLMs often introduce errors in the code. Providing opportunities for trial and error, along with error feedback mechanisms, is essential for sustained optimization. Without these tools, Gödel Agent would struggle to operate until satisfactory results are achieved. On the other hand, the code running and LLM calling have minimal impact on the outcomes, as Gödel Agent can implement these basic functionalities independently. Their inclusion at the outset primarily serves efficiency purposes.
@@ 5.2 Robustness Analysis of the Agent
Gödel Agent occasionally makes erroneous modifications, sometimes causing the agent to terminate unexpectedly or leading to degraded task performance.
Table 3: Robustness metric for Gödel Agent. Frequency of unexpected events on MGSM using CoT as the initial method.
Event Frequency (%)
Accidental Termination 4
Temporary Drop 92
Optimization Failure 14
Table 3 shows the proportion of runs on MGSM where the agent terminated, experienced performance degradation during optimization, or ultimately performed worse than its initial performance. These statistics are collected over 100 optimization trials. Thanks to the design of our error-handling mechanism, only a few percentages of agent runs result in termination. This typically occurs when Gödel Agent modifies its recursive improvement module, rendering it unable to continue self-optimization. Additionally, Gödel Agent frequently makes suboptimal modifications during each optimization iteration. However, in most cases, the final task performance surpasses the initial baseline. This indicates that Gödel Agent is able to adjust its optimization direction or revert to a previous optimal algorithm when performance declines, demonstrating the robustness in its self-improvement process.
@@ 5.3 Case Study: Game of 24
LLM Algorithms with Code-Assisted Verification:  In several runs, Gödel Agent continues to refine its LLM-based algorithm. Figure 4.a shows the improvement process, where the most significant gains come from integrating a code-assisted verification mechanism into the task algorithm and reattempting the task with additional experiential data. The former increases performance by over 10%, while the latter boosts it by more than 15%. Furthermore, Gödel Agent enhances its optimization process by not only retrieving error messages but also using the errortrace library for more detailed analysis. It adds parallel optimization capabilities, improves log outputs, and removes redundant code. These iterative enhancements in both the task and optimization algorithms show Gödel Agent ’s unique ability to continually refine itself for better performance.
@@
To analyze the impact of different initial policies on the effectiveness and efficiency of the optimization process, various methods with different levels of sophistication are used as the initial policies for the Game of 24, including Tree of Thought (ToT) (Yao et al., 2023), Chain of Thought (CoT) (Wei et al., 2022), basic prompt instructions, and prompts that deliberately produce outputs in incorrect formats not aligned with the task requirements. The results are shown in Figure 4.b.
@@
The findings indicate that stronger initial policies lead to faster convergence, with smaller optimization margins, as Gödel Agent reaches its performance limit without further enhancing its optimization capabilities. Conversely, weaker seed methods result in slower convergence and larger optimization gains, with Gödel Agent making more modifications. However, even in these cases, Gödel Agent does not outperform the results achieved using ToT. This suggests that, given the current limitations of LLMs, it is challenging for Gödel Agent to innovate beyond state-of-the-art algorithms. Improvements in LLM capabilities are anticipated to unlock more innovative self-optimization strategies in the future.
6 Discussions and Future Directions
6.1 Future Directions
6.2 Limitations
As the first self-referential agent, Gödel Agent has to construct all task-related code autonomously, which poses significant challenges. Consequently, this work does not compare directly with the most complex existing agent systems, such as OpenDevin (Wang et al., 2024b), which have benefited from extensive manual engineering efforts. Currently, Gödel Agent is not sufficiently stable and may be prone to error accumulation, hindering its ability to continue self-optimization. A more robust and advanced implementation of the Gödel Agent is anticipated, with numerous potential improvements outlined in Section 6.1. The experiments presented in this paper are intended to demonstrate the effectiveness and feasibility of recursive self-improvement.
7 Related Work
Hand-Designed Agent Systems  Researchers have designed numerous agent systems tailored to various tasks based on predefined heuristics and prior knowledge. These systems often employ techniques such as prompt engineering (Chen et al., 2023a; Schulhoff et al., 2024), chain-of-thought reasoning and planning (Wei et al., 2022; Yao et al., 2022), as well as reflection (Shinn et al., 2024; Madaan et al., 2024), code generation (Wang et al., 2023a; Vemprala et al., 2024), tool use (Nakano et al., 2021; Qu et al., 2024), retrieval-augmented generation (Lewis et al., 2020; Zhang et al., 2024b), multi-agent collaboration (Xu et al., 2023; Wu et al., 2023; Qian et al., 2023; Hong et al., 2023), and composite engineering applications (Significant Gravitas, ; Wang et al., 2024b). Once crafted by human designers, these systems remain static and do not adapt or evolve over time.
Meta-Learning Optimized Agent Systems  Some researchers have explored methods for enhancing agents through fixed learning algorithms. For example, certain frameworks store an agent’s successful or unsuccessful strategies in memory based on environmental feedback (Liu et al., 2023; Hu et al., 2023; Qian et al., 2024), while others automatically optimize agent prompts (Khattab et al., 2023; Zhang et al., 2024a; Khattab et al., 2023). Some studies have focused on designing prompts that enable agents to autonomously refine specific functions (Zhang et al., ). Zhou et al. (2024) proposed a symbolic learning framework that uses natural language gradients to optimize the structure of agents. Hu et al. (2024) used a basic foundational agent to design agents for downstream tasks. However, these algorithms for enhancement are also designed manually and remain unchanged once deployed, limiting the agents’ ability to adapt further.
Recursive Self-Improvement  The concept of recursive self-improvement has a long history (Good, 1966; Schmidhuber, 1987). Gödel machine (Schmidhuber, 2003) introduced the notion of a proof searcher that executes a self-modification only if it can prove that the modification is optimal, thereby enabling the machine to enhance itself continuously. Subsequent works by Nivel et al. (2013) and Steunebrink et al. (2016) proposed restrictive modifications to ensure safety during the self-improvement process. In the early days, there were also some discussions of self-improving agents that were not based on LLM (Hall, 2007; Steunebrink & Schmidhuber, 2012). More recently, Zelikman et al. (2023) applied recursive self-improvement to code generation, where the target of improvement was the optimizer itself, and the utility was evaluated based on performance in downstream tasks. Our proposed Gödel Agent represents the first self-improving agent where the utility function is autonomously determined by LLMs. This approach is more flexible, removing human-designed constraints and allowing the agent’s capabilities to be limited only by the foundational model itself, rather than by human design bottlenecks.
@@ 8 Conclusion
We propose Gödel Agent, a self-referential framework that enables agents to recursively improve themselves, overcoming the limitations of hand-designed agents and meta-learning optimized agents. Gödel Agent can dynamically modify its own logic based on high-level objectives. Experimental results demonstrate its superior performance, efficiency, and adaptability compared to traditional agents. This research lays the groundwork for a new paradigm in autonomous agent development, where LLMs, rather than human-designed constraints, define the capabilities of AI systems. Realizing this vision will require the collective efforts of the entire research community.
@@ C.1 Codes of the Best Policies Found by Gödel Agent Across Four Tasks
Code 1: Code of the best policy found by Gödel Agent for DROP.
@@
17 return_dict_keys=category[’return_keys’],
@@
25 if isinstance(response, list):
26 all_responses.extend(response)
27 else:
@@
54 result_solution = {key: response.get(key, []) if isinstance(response.get(key, []), list) else [response.get(key, [])] for key in final_response.keys()}
55 for key in step_counter.keys():
56 if key in result_solution.keys() and step_counter[key] and result_solution[key]:
57 final_response[’answer’] = response[’answer’]
58 final_response = result_solution
59 break
@@
Code 2: Code of the best policy found by Gödel Agent for MGSM.
@@
Code 3: Code of the best policy found by Gödel Agent for MMLU.
@@
Code 4: Code of the best policy found by Gödel Agent for GPQA.
@@ C.2 Codes in Game of 24 Tasks
Code 5: Initial code based on Chain-of-Thought for Game of 24.
@@
Code 6: Final code based on search algorithm for Game of 24.

### v3

@@ 5.3 Experimental Results and Analysis
The experimental results are shown in Table 1. Under the same setting, Gödel Agent achieves either optimal or comparable results to Meta Agent Search across all tasks. Notably, in the mathematics task MGSM, Gödel Agent outperforms it by 11%. This suggests that reasoning tasks offer greater room for improvement for Gödel Agent (performance). In contrast to Meta Agent Search, which needs to design different modules for different tasks, Gödel Agent demonstrates greater flexibility. It requires only a simple initial policy, such as CoT, with all other components being autonomously generated. Moreover, through interaction with the environment, it gradually adapts and independently devises effective methods for the current task. The final policies generated by Gödel Agent are shown in Appendix C.1. Additionally, our method converges faster, with the required number of iterations and computational cost compared to the Meta Agent shown in Appendix D.
We also conduct experiments without restrictions, where Gödel Agent significantly outperforms all baselines. Upon further analysis, we find that this is primarily due to the agent’s spontaneous requests for assistance from more powerful models such as GPT-4o in some tasks. Therefore, Gödel Agent is particularly well-suited for open-ended scenarios, where it can employ various strategies to enhance performance (potential).
Therefore, we can find that Gödel Agent is superior to the previous agent frameworks in terms of performance, flexibility, cost, and potential.
@@ 6 Analysis
To further explore how Gödel Agent self-improves, as well as its efficiency and the factors that influence it, we first evaluate the tool usage ratio on MGSM and conduct an ablation study on the initial tools. In addition, to analyze the robustness of Gödel Agent’s self-improvement, we also collect statistics for the agent’s termination. Finally, we perform a case study of initial policies and optimization processes on the classic Game of 24.
6.1 Analysis of Initial Tools
@@
We record the number of different actions taken in experiments. In Figure 4, we can see that Gödel Agent interacts with its environment frequently, analyzing and modifying its logic in the process. Additionally, error handling plays a crucial role.
Ablation MGSM Ablation MGSM
w/o think 50.8↓13.4 w/o run 57.1↓-7.1
w/o err 49.4↓-14.8 w/o LLM 60.4↓-3.8
Table 2: Ablation study on initial tool configuration. ”think” refers to ”thinking”, ”err” to ”error handling”, ”run” to ”code running”, and ”LLM” to ”LLM calling”.
As discussed in Section 4.2, Gödel Agent is initially provided with four additional tools. To analyze their impact, an ablation study is conducted, and the results are shown in Table 2. The study reveals that the “thinking before acting” tool significantly influences the results, as much of Gödel Agent’s optimization effectiveness stems from pre-action planning and reasoning. Additionally, error handling is crucial for recursive improvement, as LLMs often introduce errors in the code. Providing opportunities for trial and error, along with error feedback mechanisms, is essential for sustained optimization. On the other hand, the code running and LLM calling have minimal impact on the outcomes, as Gödel Agent can implement these basic functionalities independently. Their inclusion at the outset primarily serves efficiency purposes.
@@ 6.2 Robustness Analysis of the Agent
We test Gödel Agent on 100 optimization trials on MGSM and find it occasionally makes erroneous changes, which can result in either terminating unexpectedly (4%) or experiencing temporary performance drops (92%) during optimization. Only in 14% of trials, optimization ultimately failed, resulting in worse performance than the initial policy.
Thanks to the design of our error-handling mechanism, unexpected terminations are rare and typically occur when Gödel Agent modifies its recursive improvement module, making further self-optimization impossible. While suboptimal modifications are frequent during individual optimization steps, the final task performance usually exceeds the initial baseline. This demonstrates that Gödel Agent can adjust its optimization direction or revert to a previous optimal algorithm when performance declines, highlighting the robustness of its self-improvement process.
@@ 6.3 Case Study: Game of 24
LLM Algorithms with Code-Assisted Verification:  In several runs, Gödel Agent continues to refine its LLM-based algorithm. Figure 5.a shows the improvement process, where the most significant gains come from the code-assisted verification mechanism and reattempting the task with additional data. The former increases performance by over 10%, while the latter boosts it by more than 15%. Furthermore, Gödel Agent enhances its optimization process by not only retrieving error messages but also using the error-trace library for more detailed analysis. It adds parallel optimization capabilities, improves log outputs, and removes redundant code. These iterative enhancements in both the task and optimization algorithms show Gödel Agent’s unique ability to continually refine itself for better performance.
@@
To analyze the impact of different initial policies on the effectiveness and efficiency of optimization, various methods are used as the initial policies for the Game of 24, including Tree of Thought (ToT) (Yao et al., 2023), Chain of Thought (CoT) (Wei et al., 2022), basic prompt instructions, and prompts that deliberately produce outputs in incorrect formats not aligned with the task requirements. The results are shown in Figure 5.b.
@@
The findings indicate that stronger initial policies lead to faster convergence, with smaller optimization margins, as Gödel Agent reaches its performance limit without further enhancing its optimization capabilities. Conversely, weaker initial methods result in slower convergence and larger gains, with Gödel Agent making more modifications. However, even in these cases, Gödel Agent does not outperform the results achieved using ToT. Given the current limitations of LLMs, it is challenging for Gödel Agent to innovate beyond state-of-the-art algorithms. Improvements in LLM capabilities are anticipated to unlock more innovative self-optimization strategies in the future. We also discuss the future directions in Appendix E.
@@ 7 Conclusion
We propose Gödel Agent, a self-referential framework that enables agents to recursively improve themselves, overcoming the limitations of hand-designed agents and meta-learning optimized agents. Gödel Agent can dynamically modify its logic based on high-level objectives. Experimental results demonstrate its superior performance, efficiency, and adaptability compared to traditional agents. This research lays the groundwork for a new paradigm in autonomous agent development, where LLMs, rather than human-designed constraints, define the capabilities of AI systems.
Limitations
As the first self-referential agent, Gödel Agent has to construct all task-related code autonomously, which poses significant challenges. Consequently, this work does not compare directly with the most complex existing agent systems, such as OpenDevin (Wang et al., 2024b), which have benefited from extensive manual engineering efforts. This makes it unrealistic to expect it to outperform systems that have taken researchers several months or even years to develop. The experiments presented in this paper are intended to demonstrate the feasibility of recursive self-improvement. A more robust and advanced implementation of the Gödel Agent is anticipated, with numerous potential improvements outlined in Appendix E.
@@ Appendix A Goal Prompt of Gödel Agent
The goal prompt of Gödel Agent is shown in Box 1. It’s worth noting that this prompt has nothing to do with the downstream tasks. It merely encourages Gödel Agent to improve itself based on the environmental feedback. The agent understands the specific tasks through the environmental feedback.
Figure 6: Accuracy progression for Gödel Agent and random sampling.
@@ C.1 Codes of the Best Policies Found by Gödel Agent Across Four Tasks
Listing 1: Code of the best policy found by Gödel Agent for DROP.
@@
50 # To demonstrate, some code has been omitted.
@@
Listing 2: Code of the best policy found by Gödel Agent for MGSM.
@@
Listing 3: Code of the best policy found by Gödel Agent for MMLU.
@@
Listing 4: Code of the best policy found by Gödel Agent for GPQA.
@@ C.2 Codes in Game of 24 Tasks
Listing 5: Initial code based on Chain-of-Thought for Game of 24.
@@
Listing 6: Final code based on search algorithm for Game of 24.
@@ Appendix D Cost of Experiments
Appendix E Discussions and Future Directions
Appendix F Additional Novel Policies Designed by Gödel Agent
In this section, we present the optimization process of Gödel Agent on MGSM, illustrating its progress across various iteration steps within a single optimization run. The strategy obtained in the 6th iteration (shown in Code 7) reflects the Gödel Agent’s comprehension of mathematical tasks, attempting to handle them through a process akin to parse-deduct-execute-validate. By the 14th iteration, as illustrated in Code 8, the strategy evolves through the summarization of erroneous cases, abstracting key insights and employing a checklist to guide the validation process. Finally, the strategy at the 20th iteration (demonstrated in Code 9) asserts the use of a ”rabbit-proof syntax tactline, reinforced by consistent effort through role-coded checks,” to refine prompt design. In the end, we also show one analysis example of Gödel Agent.
Appendix G Comparison Between Random Sampling and Gödel Agent Performance
To demonstrate the distinction between our approach and random sampling, we conducted 30 independent random sampling experiments using GPT-4o. The prompts used for random sampling were identical to the initial policy prompts employed by Gödel Agent to ensure a fair comparison. The results are illustrated in Figure 6. From the figure, it is evident that the performance of random sampling remains around 30% across all trials. In contrast, Gödel Agent, despite experiencing occasional temporary dips in performance, rapidly corrects these deviations and demonstrates continuous improvement over iterations. This consistent upward trajectory highlights the superiority of Gödel Agent over random sampling. The Gödel Agent’s ability to leverage feedback and recursively optimize its policies underscores its effectiveness in achieving higher performance.
Listing 7: Policy at 6th Iteration found by Gödel Agent for MGSM.
1def solver(agent, task: str):
2 def parse_problem(task):
3 # Basic arithmetic and logical parsing based on keywords
4 words = task.split()
5 numbers = list(map(int, filter(lambda x: x.isdigit(), words)))
6 return {’numbers’: numbers, ’text’: task}
7
8 def perform_logic_deduction(parsed_details):
9 # make deductions based on common problem formats
10 numbers = parsed_details[’numbers’]
11 # This will only manage simple sum, subtraction, multiplication inference
12 logic_map = {
13 ’add’: lambda a, b: a + b,
14 ’subtract’: lambda a, b: a - b,
15 ’multiply’: lambda a, b: a * b
16 }
17 # Try to identify actions based on keywords
18 if ’sum’ in parsed_details[’text’] or ’total’ in parsed_details[’text’]:
19 result = sum(numbers)
20 elif ’difference’ in parsed_details[’text’] or ’less’ in parsed_details[’text’]:
21 result = logic_map[’subtract’](numbers[0], numbers[1])
22 elif ’product’ in parsed_details[’text’] or ’times’ in parsed_details[’text’]:
23 result = logic_map[’multiply’](numbers[0], numbers[1])
24 else:
25 # Default case showing no deduction
26 result = 0
27 return result
28
29 def execute_computation(logic_results):
30 # Taking result from inference to numerical handling
31 return logic_results
32
33 def validate_and_compile_results(computation_results):
34 # Prepares and ensures the response matches expected format
35 final_answer = computation_results
36 return final_answer
37
38 try:
39 # Parsing
40 parsed_details = parse_problem(task)
41
42 # Logical deduction
43 logic_results = perform_logic_deduction(parsed_details)
44
45 # Computation
46 computation_results = execute_computation(logic_results)
47
48 # Validation and compilation
49 final_answer = validate_and_compile_results(computation_results)
50
51 return {"answer": final_answer}
52 except Exception as e:
53 return {"error": str(e)}
Listing 8: Policy at 14th Iteration found by Gödel Agent for MGSM.
1def solver(agent, task: str):
2 # Updated examples to mirror tasks needing layered logical verification.
3 examples = [
4 {’description’: ’Task exemplifying standard integer-based calculations.’, ’reasoning’: ’Use arithmetic transformations to validate expressions and correct errors if any arise, ensuring correctness.’, ’answer’: 20},
5 {’description’: ’Example to validate word problem conversion to math.’, ’reasoning’: ’Stepwise interpretation from words into math operations and bridge which logic errors need capture.’, ’answer’: 15},
6 {’description’: ’Scenario involving normalizing uneven division instances.’, ’reasoning’: ’Ensure no division by zero and equal verification of logical conclusions.’, ’answer’: 6},
7 ]
8
9 # Task prompt incorporating roles with enhanced checklists after operation conclusion.
[…]
13 task_prompt += ’; .join([f"{ex[’description’]} -> Reasoning: {ex[’reasoning’]} | Answer: {ex[’answer’]}" for ex in examples])
[…]
15
16 messages = [{"role": "user", "content": task_prompt}]
17
18 response = agent.action_call_json_format_llm(
19 model="gpt-3.5-turbo",
20 messages=messages,
21 temperature=0.3,
22 num_of_response=1,
23 role="solution auditor",
24 return_dict_keys=["description", "reasoning", "answer"],
25 requirements=(
26 "1. Validate arithmetic consistency and integrity within calculations."
27 "2. Utilize any corrections to refine answer outputs incrementally."
28 ).strip(),
29 )
30
31 return_dict = response[0]
32 return_dict["answer"] = str(return_dict.get("answer", ""))
33 return return_dict
Listing 9: Policy at 20th Iteration found by Gödel Agent for MGSM.
1def solver(agent, task: str):
2 # Targets design for specific error-prone areas with preceding misfires.
3 examples = [
4 {’description’: ’Immediate Arithmetic Operations’, ’reasoning’: ’Observe step-by-step through a chain of logical confirmations.’, ’answer’: 20},
5 {’description’: ’Sequential Word Problem Breakdown’, ’reasoning’: ’Ensure smaller module segment steps match logical math outputs consistently.’, ’answer’: 15},
6 {’description’: ’Fraction and Cascade Operations’, ’reasoning’: ’Validate each fraction conversion before proceeding to other forms.’, ’answer’: 6},
7 ]
8
9 # Building a rabbit-proof syntax tactline, pushing consistent effort from role-coded checks.
[…]
13 task_prompt += ’; .join([f"{ex[’description’]} -> Rationale: {ex[’reasoning’]} | Response: {ex[’answer’]}" for ex in examples])
[…]
15
16 messages = [{"role": "user", "content": task_prompt}]
17
18 response = agent.action_call_json_format_llm(
19 model="gpt-3.5-turbo",
20 messages=messages,
21 temperature=0.3,
22 num_of_response=1,
23 role="logical verifier",
24 return_dict_keys=["description", "reasoning", "answer"],
25 requirements=(
26 "1. Review arithmetic sessions for comprehensive error fixation."
27 "2. Develop clarity with resistant focus on authentic semantic directive flows."
28 ).strip(),
29 )
30
31 return_dict = response[0]
32 return_dict["answer"] = str(return_dict.get("answer", ""))
33 return return_dict
Analysis Example of Gödel Agent The current logic of the solver in the agent_module is achieving only a 45% accuracy on the validation set. The structure of the solver involves a single call to the LLM with a specification that the answer must be an integer. While this approach seems correct, it may not be sufficient to tackle more complex problems where multipart reasoning or verification is needed, leading to significant errors. Problems Identified: 1. Integer Conversion Misunderstandings: Some math problems aren’t resulting in integer answers natively, affecting accuracy when the solver forces an integer format. 2. Reasoning Oversights: The LLM, designated as ”math expert,” occasionally misses necessary reasoning steps or executes incorrect operations leading to incorrect answers (like sample 10). 3. Ambiguity in LLM Outputs: Given tasks in multiple languages, there might be ambiguity in math translations or formulations affecting comprehension (like sample 6). Plan for Improvement: 1. Elaborate Reasoning: Consider a multi-step LLM call to allow for a reasoning review before finalizing the answer. 2. Intelligent Answer Parsing: Instead of forcibly converting to integers, parse the LLM response intelligently to see if the answer justifies integer rounding or alterations. 3. Diverse Language Handling: Create a pre-processing step to consistently translate non-English math problems into English for uniform processing. 4. Examples & Feedback: Utilize correctly solved examples as context if a single-step LLM fails. Next Steps: 1. Enhance Solver Logic: Implement a multi-step process involving interim LLM validation or rationale cross-checking. 2. Pre-process Language Variations: Ensure all non-English tasks are translated and re-assessed. 3. Include Example Feedback: Use correctly solved examples from the evaluation as references to guide LLM reasoning and answer validation.